# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 3 Coderivative-Based Newton Method for Subgradient Systems
when $k$ is sufficiently large. This means that the quotient $\|x^{k+1}-\bar{x}\|/\|x^{k}-\bar{x}\|$ converges to 00 as $k\to\infty$, which thus completes the proof of the theorem.
Remark 3.12 (on the proof of Theorem 3.11 and comparison with inexact generalized Newton methods).
Following the referee’s suggestion, we present an alternative proof of Theorem 3.11 by modifying the convergence results of the inexact Newton method developed in [33].
From Proposition 2.5, due to the continuous prox-regularity of $\varphi$ at $\bar{x}$ for 00, the function $e_{\lambda}\varphi$ is of class $\mathcal{C}^{1,1}$ on a neighborhood $U_{\lambda}$ of $\bar{x}$. The referee suggests modifying the inexact Newton method from [33, Theorem 3] to solve the equation $F(z)=0$, where $F(z)=\nabla e_{\lambda}\varphi(z)$. The sequence $\{z^{k}\}$ is constructed as follows: Choose $z^{k}\in U_{\lambda}$ and find $p^{k}\in{\rm I\!R}^{n}$ satisfying the conditions
$\text{\rm dist}(p^{k},[D^{*}F(z^{k})]^{-1}(-F(z^{k}))\leq\alpha_{k}\|z^{k}-\bar{x}\|\quad\text{and }\;z^{k+1}=z^{k}+p^{k},\;k\in{\rm I\!N}.$ (3.24)
In the approximation step of Algorithm 1, we have $\widehat{v}^{k}\in\partial\varphi(\widehat{x}^{k})$ for which (3.10) holds. Define
$z^{k}:=\widehat{x}^{k}+\lambda\widehat{v}^{k},\quad p^{k}:=z^{k+1}-z^{k},$ (3.25)
which leads us to $z^{k}=\widehat{x}^{k}+\lambda\nabla e_{\lambda}\varphi(z^{k})$. Then Lemma 3.9 yields
$d^{k}-\lambda\nabla e_{\lambda}\varphi(z^{k})\in[D^{*}F(z^{k})]^{-1}(-F(z^{k})),$
and thus we arrive at the relationships
$\text{\rm dist}(p^{k},[D^{*}F(z^{k})]^{-1}(-F(z^{k}))\leq\left\|p_{k}-d^{k}+\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right\|=\left\|z^{k+1}-z^{k}-d^{k}+\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right\|$
$=\left\|z^{k+1}-\widehat{x}^{k+1}+\left(\widehat{x}^{k+1}-x^{k+1}\right)+\left(x^{k+1}-z^{k}\right)-d^{k}+\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right\|$
$=\left\|\lambda\nabla e_{\lambda}\varphi\left(z^{k+1}\right)+\left(\widehat{x}^{k+1}-x^{k+1}\right)+\left(x^{k+1}-\widehat{x}^{k}-\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right)-d^{k}+\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right\|$
$=\left\|\lambda\nabla e_{\lambda}\varphi\left(z^{k+1}\right)+\left(\widehat{x}^{k+1}-x^{k+1}\right)+\left(x^{k+1}-\widehat{x}^{k}\right)-d^{k}\right\|$
$=\left\|\lambda\hat{v}^{k+1}+\left(\widehat{x}^{k+1}-\bar{x}\right)-\left(x^{k+1}-\bar{x}\right)\right\|\leq\left\|\lambda\hat{v}^{k+1}+\left(\widehat{x}^{k+1}-\bar{x}\right)\right\|+\left\|\left(x^{k+1}-\bar{x}\right)\right\|$
$\leq\sqrt{2}\left\|\left(\widehat{x}^{k+1},\widehat{v}^{k+1}\right)-(\bar{x},0)\right\|+\left\|x^{k+1}-\bar{x}\right\|\leq(\sqrt{2}\eta+1)\left\|x^{k+1}-\bar{x}\right\|.$
Using the semismoothness of $F$ and proceeding as in the proof of [33, Theorem 3] allow us to get the following assertions provided that $x^{0}$ is sufficiently close to $\bar{x}$:
$\bullet$ The sequences $\left\{x^{k}\right\}$ and $\left\{z^{k}\right\}$ converge to $\bar{x}$.
$\bullet$ $\left\|z^{k+1}-\bar{x}\right\|=o\left(\left\|z^{k}-\bar{x}\right\|\right)=o\left(\left\|x^{k}-\bar{x}\right\|\right)$.
$\bullet$ $\left\|x^{k+1}-\bar{x}\right\|=\left\|\widehat{x}^{k}+d^{k}-\bar{x}\right\|=\left\|z^{k}+\left(d^{k}-\lambda\nabla e_{\lambda}\varphi\left(z^{k}\right)\right)-\bar{x}\right\|=o\left(\left\|z^{k}-\bar{x}\right\|\right)=o\left(\left\|x^{k}-\bar{x}\right\|\right)$.
$\bullet$ The sequence $\left\{\alpha_{k}:=\displaystyle\frac{(\sqrt{2}\eta+1)\left\|x^{k+1}-\bar{x}\right\|}{\left\|z^{k}-\bar{x}\right\|}\right\}$ converges to 0.
While we gratefully acknowledge the proof advised by the referee, it seems to us that this proof lacks natural simplicity. The suggested approach is not straightforward since it requires modifications of the algorithm in [33] and additional inequality evaluations. In contrast, our proof is more direct and avoids excessive auxiliary results. Moreover, Algorithm 1 is more practical than the modified inexact Newton method in (3.24). Let us present the detailed comparisons regarding practical aspects:
$\bullet$ Initial points. The implementation of Algorithm (3.24) suggested by the referee requires that the initial point $z^{0}$ lies within $U_{\lambda}$, where the Moreau envelope $e_{\lambda}\varphi$ is of class $\mathcal{C}^{1,1}$. In contrast, Algorithm 1 does not impose such a requirement since it does not depend on the single-valuedness of the proximal mapping except that near the solution $\bar{x}$. This provides more flexibility for globalization strategies as discussed in Section 6 below.
$\bullet$ Computational aspects. The sequence generated by (3.24) relies heavily on the explicit formula for the Moreau envelope $e_{\lambda}\varphi$, which is often difficult to compute, particularly when $\varphi$ is either a composite function, or it is the sum of two nonconvex functions. As such, this method is more valuable for qualitative analysis than for quantitative computations. In contrast, Algorithm 1 involves the computation of first- and second-order subdifferentials at each step without requiring the formulation of Moreau envelopes. This approach takes advantage of well-established subdifferential calculus, and—for many practical classes of extended-real-valued functions—the explicit computations of the second-order subdifferentials are widely accessible; see, e.g., [41].
$\bullet$ Applicable classes of problems. Although solving $0\in\partial\varphi(x)$ may be transformed in some settings into solving a locally Lipschitz equation $F(z)=0$, such a reformulation is not always straightforward. Moreover, computing $F$ iteratively may be infeasible when it requires the explicit form of the Moreau envelope. On the other hand, Algorithm 1 offers a direct approach to solving nonsmooth and nonconvex optimization problems without the need for such reformulations. This generality enables our method to handle problems involving the sum of two nonconvex functions as detailed in the remainder of this paper. Observe, e.g., that for the classes of optimization problems considered in Section 7, reformulating them as equivalent locally Lipschitz equations is particularly challenging due to the intrinsic nonconvexity of the $\ell_{0}$ norm. Moreover, our approach lays a solid foundation for future research on broader classes of nonconvex optimization problems including composite optimization, difference-of-convex programming, and related models. Let us emphasize that subdifferential calculus and computations for many of these structures are already well developed.
@@
(i) The semismooth Newton method of [13] addresses solving the inclusion $0\in F(x)$ for $F:{\rm I\!R}^{n}\rightrightarrows{\rm I\!R}^{n}$, which reduces to (3.1) when $F:=\partial\varphi$ for some $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$. The underlying difference between our approach and the method of [13] is in finding generalized Newton directions. To be more specific, the directions $d^{k}$ in the semismooth Newton method is given by
$d^{k}:=-A^{-1}B\widehat{v}^{k},\quad k=0,1,\ldots,$ (3.26)
@@
(ii) Our previous paper [24] develops local coderivative-based Newton methods to solve inclusions of the form $0\in\partial\varphi(x)$. The algorithms proposed therein heavily rely on the proximal mapping of $\varphi$, which is not always easy to compute. Moreover, in [24] we focus on the case where $\varphi=f+g$, with $f$ being quadratic and $g$ prox-regular. In contrast, our new Algorithm 1 does not require such structural assumptions. We consider a more general setting where $\varphi$ may be nonsmooth and nonconvex, and our method only requires the computation of first- and second-order subdifferentials, which enjoy full calculi and can often be determined explicitly.
@@ 4 Local Convergence Analysis of Coderivative-Based Newton Method for Structured Nonconvex Optimization
Let $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be given in (4.1). For each $\lambda\in(0,\lambda_{g})$, define the following multifunctions:
@@
find $\text{find }\bar{x}\in\Omega\;\text{ such that }\;\langle F(\bar{x}),x-\bar{x}\rangle\geq 0\;\mbox{ for all }\;x\in\Omega,$ (4.5)
@@
Note that both normal maps in (4.3) and (4.6) agree when $F:=\nabla f$, $g:=\delta_{\Omega}$, and $\lambda=1$. Various error bounds for variational inequalities based on the normal map (4.6) can be found in the book [11] and the references therein, while new ones by using this normal map have been recently obtained in [20] for the class of strongly pseudomonotone variational inequalities. One of the first extensions of the normal map to composite problems was introduced by Pieper in his PhD thesis [47]. His approach combines semismooth Newton steps for solving the nonsmooth equation $F^{\text{nor}}_{\lambda}(x)=0$ with a trust-region-type globalization mechanism. The subsequent paper [32] analyzed local properties of a quasi-Newton variant of Pieper’s normal map-based trust-region method. Pieper’s approach has also been successfully applied in various contexts, particularly in optimal control; see, e.g., [6, 31, 56]. A more comprehensive convergence analysis, covering both local and global aspects of the normal map trust-region method, was recently presented in [44]. Additionally, algorithmic extensions of normal map-based techniques to stochastic settings have been proposed in [34, 50]. In [44], the authors introduce the normal map as an intermediate step to find a solution to the problem of minimizing the sum of a smooth and a nonsmooth convex functions by applying the semismooth Newton method to the equation $\mathcal{F}_{\lambda}^{\text{\rm nor}}(z)=0$, where the local Lipschitz continuity of this mapping is ensured by the convexity of the regularizer. In contrast, our approach bypasses this transformation and directly tackles the optimization problem, where the objective function is the sum of two nonconvex functions. In this framework, the normal map plays a critical role in implementing the approximate step of Algorithm 1 and in supporting the development of our new generalized Newton and line-search methods introduced in this section and further detailed in Sections 5 and 6.
@@
Let $\varphi$ be given in (4.1), and let $\bar{x}\in\mbox{\rm dom}\,\varphi$ Then $\bar{x}$ is:
(i) F((((réchet))))-stationary point if $0\in\widehat{\partial}\varphi(\bar{x})$.
(ii) M((((ordukhovich))))-stationary point if $0\in{\partial}\varphi(\bar{x})$.
(iii) Critical point if $\bar{x}\in\mathcal{G}_{\lambda}(\bar{x})$ for some $\lambda\in(0,\lambda_{g})$.
@@
If $\bar{x}$ is a critical point of (4.1), then $\bar{x}$ is a F-stationary (and hence M-stationary) point. Indeed, we have from the critical point definition that
@@
It then follows from [39, Proposition 1.114] that
@@
If $f$ and $g$ are convex and $\bar{x}$ is a critical point of (4.1), then $\bar{x}$ is an optimal solution to (4.1); see [4, Corollary 27.3].
@@
Let $\varphi$ be given in (4.1), and let $\bar{x}$ be an optimal solution to problem (4.1). Assume that $\nabla f$ is Lipschitz continuous on ${\rm I\!R}^{n}$ with modulus $L_{f}>0$. Then $\bar{x}$ is a critical point of (4.1), and $\mathcal{G}_{\lambda}$ is single-valued at $\bar{x}$ for any $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$.
@@
Since $\bar{x}$ is an optimal solution to (4.1) and $g$ is proper, it tells us that $g(\bar{x})$ is finite, and so is $g(\widehat{x})$ by (4.8). Combining (4.7) and (4.8) provides the conditions
$f(\widehat{x})+g(\widehat{x})$ $\leq f(\bar{x})+\langle\nabla f(\bar{x}),\widehat{x}-\bar{x}\rangle+\frac{L_{f}}{2}\|\widehat{x}-\bar{x}\|^{2}+g(\widehat{x})$
@@
[…]
@@
Let $\varphi$ be given in (4.1). Then for each $\lambda\in(0,\lambda_{g})$, the following assertions hold:
(i) If $\bar{x}$ is a critical point of (4.1), then $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$.
(ii) If $\nabla f$ is Lipschitz continuous around $\bar{x}$ with modulus $L_{f}>0$ and if $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$ with $\lambda\in(0,1/L_{f})$, then $\bar{x}$ is a critical point of (4.1).
[Proof.] To verify (i), we get by the criticality $\bar{x}$ in (4.1) that $\bar{x}\in\mathcal{G}_{\lambda}(\bar{x})$, which yields
@@
$0=\nabla f(\widehat{x})+\frac{1}{\lambda}\big{(}\bar{x}-\lambda\nabla f(\bar{x})-\widehat{x}\big{)}=\nabla f(\widehat{x})-\nabla f(\bar{x})+\frac{1}{\lambda}(\bar{x}-\widehat{x}).$ (4.9)
@@
$\frac{1}{\lambda}\|\widehat{x}-\bar{x}\|=\|\nabla f(\widehat{x})-\nabla f(\bar{x})\|\leq L_{f}\|\widehat{x}-\bar{x}\|,$ (4.10)
@@
Let $\varphi$ be given in (4.1). For each $\lambda\in(0,\lambda_{g})$ and $x\in{\rm I\!R}^{n}$, take $\widehat{x}\in\mathcal{G}_{\lambda}(x)$ and denote
$\widehat{v}:=\nabla f(\widehat{x})-\nabla f(x)+\frac{1}{\lambda}(x-\widehat{x}).$
@@
It follows from the elementary calculus rules for regular subgradients [40, Proposition 1.30] that
@@
$\widehat{v}=\nabla f(\widehat{x})-\nabla f(x)+\frac{1}{\lambda}(x-\widehat{x})\in\nabla f(\widehat{x})+\widehat{\partial}g(\widehat{x})=\widehat{\partial}\varphi(\widehat{x}),$
@@
Let $\bar{x}$ be a critical point of (4.1), and let $g$ be prox-regular with modulus $r>0$ at $\bar{x}$ for $-\nabla f(\bar{x})$. Then for each $\lambda\in(0,r^{-1})$, there exists a neighborhood $U$ of $\bar{x}$ on which the maps $x\mapsto\mathcal{G}_{\lambda}(x)$ and $x\mapsto\mathcal{F}_{\lambda}^{\text{\rm nor}}(x-\lambda\nabla f(x))$ are single-valued and Lipschitz continuous on $U$. Moreover, we find $\eta_{1},\eta_{2}>0$ such that
@@
[Proof.] Since $f$ is $\mathcal{C}^{2}$-smooth, there is a neighborhood $\widetilde{U}$ of $\bar{x}$ such that $\nabla f$ is Lipschitz continuous on $\widetilde{U}$. It follows from Lemma 2.5 that the proximal mapping $\text{\rm Prox}_{\lambda g}$ is single-valued and locally Lipschitzian on a neighborhood $U_{\lambda}$ of $\bar{x}-\lambda\nabla f(\bar{x})$ and that $\text{\rm Prox}_{\lambda g}(\bar{x}-\lambda\nabla f(\bar{x}))=\bar{x}$. Choose a neighborhood $U\subset\widetilde{U}$ of $\bar{x}$ so that
@@
$\mathcal{F}_{\lambda}^{\text{\rm nor}}\big{(}x-\lambda\nabla f(x)\big{)}=\nabla f\big{(}\mathcal{G}_{\lambda}(x)\big{)}-\nabla f(x)+\frac{1}{\lambda}\big{(}x-\mathcal{G}_{\lambda}(x)\big{)}\;\mbox{ for all }\;x\in U,$ (4.13)
@@
In the setting of Proposition 4.7, for each $\lambda\in(0,r^{-1})$ there exist $\eta>0$ and a neighborhood $U$ of $\bar{x}$ such that whenever $x\in U$, we get a pair $(\widehat{x},\widehat{v})\in\mbox{\rm gph}\,\partial\varphi$ satisfying the estimate
@@
$\widehat{x}:=\mathcal{G}_{\lambda}(x)\;\text{ and }\;\widehat{v}:=\mathcal{F}_{\lambda}^{\text{\rm nor}}(x-\lambda\nabla f(x)).$
@@
$\widehat{x}^{k}\in\text{\rm Prox}_{\lambda g}(x^{k}-\lambda\nabla f(x^{k})),\quad\widehat{v}^{k}:=\nabla f(\widehat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\widehat{x}^{k})$ (4.15)
@@
$-\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\in\partial^{2}g(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k})(d^{k})$ (4.16)
@@
$-\widehat{v}^{k}\in\nabla^{2}f(\widehat{x}^{k})d^{k}+\partial^{2}g\big{(}\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k})\big{)}(d^{k}),$
which is equivalent to $-\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$ by the second-order sum rule from [39, Proposition 1.121].

### v3

@@ 3 Coderivative-Based Newton Method for Subgradient Systems
when $k$ is sufficiently large. This means that the quotient $\|x^{k+1}-\bar{x}\|/\|x^{k}-\bar{x}\|$ converges to 0 as $k\to\infty$, which thus completes the proof of the theorem.
@@
(i) The semismooth Newton method of [13] addresses solving the inclusion $0\in F(x)$ for $F:{\rm I\!R}^{n}\rightrightarrows{\rm I\!R}^{n}$, which reduces to (3.1) when $F:=\partial\varphi$ for some $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$. The underlying difference between our approach and the method of [13] is in finding generalized Newton directions. To be more specific, the directions $d^{k}$ in the semismooth Newton method are given by
$d^{k}:=-A^{-1}B\widehat{v}^{k},\quad k=0,1,\ldots,$
@@
(ii) Our previous paper [24] develops local coderivative-based Newton methods to solve inclusions of the form $0\in\partial\varphi(x)$. The algorithms proposed therein heavily rely on the proximal mapping of $\varphi$, which is not always easy to compute. Moreover, in [24] we focus on the case where $\varphi=f+g$, with $f$ being quadratic and $g$ prox-regular. In contrast, our new Algorithm 1 does not require such structural assumptions.
@@ 4 Local Convergence Analysis of Coderivative-Based Newton Method for Structured Nonconvex Optimization
Let $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be given as in (4.1). For each $\lambda\in(0,\lambda_{g})$, define the following multifunctions:
@@
$\text{find }\bar{x}\in\Omega\;\text{ such that }\;\langle F(\bar{x}),x-\bar{x}\rangle\geq 0\;\mbox{ for all }\;x\in\Omega,$ (4.5)
@@
Note that both normal maps in (4.3) and (4.6) agree when $F:=\nabla f$, $g:=\delta_{\Omega}$, and $\lambda=1$. Various error bounds for variational inequalities based on the normal map (4.6) can be found in the book [11] and the references therein, while new ones by using this normal map have been recently obtained in [20] for the class of strongly pseudomonotone variational inequalities. One of the first extensions of the normal map to composite problems was introduced by Pieper in his PhD thesis [46]. His approach combines semismooth Newton steps for solving the nonsmooth equation $\mathcal{F}^{\text{nor}}_{\lambda}(x)=0$ with a trust-region-type globalization mechanism. The subsequent paper [32] analyzed local properties of a quasi-Newton variant of Pieper’s normal map-based trust-region method. Pieper’s approach has also been successfully applied in various contexts, particularly in optimal control; see, e.g., [6, 31, 55]. A more comprehensive convergence analysis, covering both local and global aspects of the normal map trust-region method, was recently presented in [43]. Additionally, algorithmic extensions of normal map-based techniques to stochastic settings have been proposed in [33, 49]. In [43], the authors introduce the normal map as an intermediate step to find a solution to the problem of minimizing the sum of a smooth and a nonsmooth convex functions by applying the semismooth Newton method to the equation $\mathcal{F}_{\lambda}^{\text{\rm nor}}(z)=0$, where the local Lipschitz continuity of this mapping is ensured by the convexity of the regularizer. In contrast, our approach bypasses this transformation and directly tackles the optimization problem, where the objective function is the sum of two nonconvex functions. In this framework, the normal map plays a critical role in implementing the approximate step of Algorithm 1 and in supporting the development of our new generalized Newton and line-search methods introduced in this section and further detailed in Sections 5 and 6.
@@
Let $\varphi$ be given as in (4.1), and let $\bar{x}\in\mbox{\rm dom}\,\varphi$. Then $\bar{x}$ is:
(i) F((réchet))-stationary point if $0\in\widehat{\partial}\varphi(\bar{x})$.
(ii) M((ordukhovich))-stationary point if $0\in{\partial}\varphi(\bar{x})$.
(iii) $\lambda$-critical point if $\bar{x}\in\mathcal{G}_{\lambda}(\bar{x})$ where $\lambda\in(0,\lambda_{g})$.
@@
Observe the following:
(i) If $\bar{x}$ is a $\lambda$-critical point of (4.1) for some $\lambda\in(0,\lambda_{g})$, then $\bar{x}$ is also a $\lambda^{\prime}$-critical point of (4.1) for any $\lambda^{\prime}\in(0,\lambda].$ Indeed, fix such $\lambda^{\prime}$, the $\lambda$-criticality of $\bar{x}$ implies that $\bar{x}\in\mathcal{G}_{\lambda}(\bar{x})$, i.e.,
$g(x)+\frac{1}{2\lambda}\|x-\bar{x}+\lambda\nabla f(\bar{x})\|^{2}\geq g(\bar{x})+\frac{1}{2\lambda}\|\bar{x}-\bar{x}+\lambda\nabla f(\bar{x})\|^{2}\quad\text{for all }\;x\in{\rm I\!R}^{n},$
which ensures in turn that
$g(x)\geq g(\bar{x})+\langle-\nabla f(\bar{x}),x-\bar{x}\rangle-\frac{1}{2\lambda}\|x-\bar{x}\|^{2}\geq g(\bar{x})+\langle-\nabla f(\bar{x}),x-\bar{x}\rangle-\frac{1}{2\lambda^{\prime}}\|x-\bar{x}\|^{2}$
for all $x\in{\rm I\!R}^{n},$ i.e., we have the estimate
$g(x)+\frac{1}{2\lambda^{\prime}}\|x-\bar{x}+\lambda^{\prime}\nabla f(\bar{x})\|^{2}\geq g(\bar{x})+\frac{1}{2\lambda^{\prime}}\|\bar{x}-\bar{x}+\lambda^{\prime}\nabla f(\bar{x})\|^{2}\quad\text{for all }\;x\in{\rm I\!R}^{n}.$
This means that $\bar{x}\in\mathcal{G}_{\lambda^{\prime}}(\bar{x})$, or $\bar{x}$ is a $\lambda^{\prime}$-critical point of (4.1).
(ii) If $\bar{x}$ is a $\lambda$-critical point of (4.1) for some $\lambda\in(0,\lambda_{g})$, then $\bar{x}$ is a F-stationary (and hence M-stationary) point. Indeed, we have from the critical point definition that
@@
It then follows from [38, Proposition 1.114] that
@@
If $f$ and $g$ are convex and $\bar{x}$ is a $\lambda$-critical point of (4.1) for some $\lambda>0$, then $\bar{x}$ is an optimal solution to (4.1); see [4, Corollary 27.3].
@@
Let $\varphi$ be given as in (4.1), and let $\bar{x}$ be an optimal solution to problem (4.1). Assume that $\nabla f$ is Lipschitz continuous on ${\rm I\!R}^{n}$ with modulus $L_{f}>0$. Then $\bar{x}$ is a $\lambda$-critical point of (4.1), and $\mathcal{G}_{\lambda}$ is single-valued at $\bar{x}$ for any $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$.
@@
Since $\bar{x}$ is an optimal solution to (4.1) and $g$ is proper, it tells us that $g(\bar{x})$ is finite, and so is g(x^)g($\widehat${x}) by (4.8). Combining (4.7) and (4.8) provides the conditions
[…]
@@
$\text{optimal solution}\Longrightarrow\text{critical point}\Longrightarrow\text{F-stationary point}\Longrightarrow\text{M-stationary point}.$
@@
Let $\varphi$ be given as in (4.1). Then for each $\lambda\in(0,\lambda_{g})$, the following assertions hold:
(i) If $\bar{x}$ is a $\lambda$-critical point of (4.1), then $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$.
(ii) If $\nabla f$ is Lipschitz continuous around $\bar{x}$ with modulus $L_{f}>0$ and if $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$ with $\lambda\in(0,1/L_{f})$, then $\bar{x}$ is a $\lambda$-critical point of (4.1).
[Proof.] To verify (i), we get by the $\lambda$-criticality $\bar{x}$ in (4.1) that $\bar{x}\in\mathcal{G}_{\lambda}(\bar{x})$, which yields
@@
[…]
@@
$\frac{1}{\lambda}\|\widehat{x}-\bar{x}\|=\|\nabla f(\widehat{x})-\nabla f(\bar{x})\|\leq L_{f}\|\widehat{x}-\bar{x}\|,$
@@
Let $\varphi$ be given as in (4.1). For each $\lambda\in(0,\lambda_{g})$ and $x\in{\rm I\!R}^{n}$, take $\widehat{x}\in\mathcal{G}_{\lambda}(x)$ and denote
[…]
@@
It follows from the elementary calculus rules for regular subgradients [39, Proposition 1.30] that
@@
[…]
@@
Let $\bar{x}$ be an M-stationary point of (4.1), and let $g$ be $r$-level prox-regular at $\bar{x}$ for $\bar{v}:=-\nabla f(\bar{x})$ for some $r>0$. Suppose that $\bar{x}$ is a $\bar{\lambda}$-critical point of (4.1) for some $\bar{\lambda}\in(0,1/r].$ Then for each $\lambda\in\left(0,\bar{\lambda}\right)$, there exists a neighborhood $U$ of $\bar{x}$ on which the maps $x\mapsto\mathcal{G}_{\lambda}(x)$ and $x\mapsto\mathcal{F}_{\lambda}^{\text{\rm nor}}(x-\lambda\nabla f(x))$ are single-valued and Lipschitz continuous on $U$. Moreover, we find $\eta_{1},\eta_{2}>0$ such that
@@
[Proof.] Since $f$ is $\mathcal{C}^{2}$-smooth, there is a neighborhood $\widetilde{U}$ of $\bar{x}$ such that $\nabla f$ is Lipschitz continuous on $\widetilde{U}$. Since $\bar{x}$ is a $\bar{\lambda}$-critical point of (4.1), then $\bar{x}\in\mathcal{G}_{\bar{\lambda}}(\bar{x})=\text{\rm Prox}_{\bar{\lambda}g}\big(\bar{x}+\bar{\lambda}\bar{v}\big)$. It follows from Proposition 2.5 and Remark 2.6(iii) that for any $\lambda\in(0,\bar{\lambda})$, the proximal mapping $\text{\rm Prox}_{\lambda g}$ is single-valued and locally Lipschitzian on a neighborhood $U_{\lambda}$ of $\bar{x}+\lambda\bar{v}$ and that $\text{\rm Prox}_{\lambda g}(\bar{x}+\lambda\bar{v})=\bar{x}$. Choose a neighborhood $U\subset\widetilde{U}$ of $\bar{x}$ so that
@@
$\mathcal{F}_{\lambda}^{\text{\rm nor}}\big(x-\lambda\nabla f(x)\big)=\nabla f\big(\mathcal{G}_{\lambda}(x)\big)-\nabla f(x)+\frac{1}{\lambda}\big(x-\mathcal{G}_{\lambda}(x)\big)\;\mbox{ for all }\;x\in U,$
@@
In the setting of Proposition 4.7, for each $\lambda\in(0,\bar{\lambda})$ there exist $\eta>0$ and a neighborhood $U$ of $\bar{x}$ such that whenever $x\in U$, we get a pair $(\widehat{x},\widehat{v})\in\mbox{\rm gph}\,\partial\varphi$ satisfying the estimate
@@
[…]
@@
$\widehat{x}^{k}\in\text{\rm Prox}_{\lambda g}(x^{k}-\lambda\nabla f(x^{k})),\quad\widehat{v}^{k}:=\nabla f(\widehat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\widehat{x}^{k})$
@@
$\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\in\partial^{2}g(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k})(d^{k})$
@@
$\widehat{v}^{k}\in\nabla^{2}f(\widehat{x}^{k})d^{k}+\partial^{2}g\big(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k})\big)(d^{k}),$
which is equivalent to $\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$ by the second-order sum rule from [38, Proposition 1.121].